OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 – 6x – 7. Find f(g(2)). :)))

OpenStudy (anonymous):

do you know how to find \[g(2)\]?

OpenStudy (anonymous):

no clue, i read through this lesson and didn't understand a thing..

OpenStudy (anonymous):

ok \[g(x)=x^2-6x-7\] so \[g(2)\] means replace x by 2

OpenStudy (anonymous):

i.e. where you see "x" you put "2" and get a number, namely \[g(2)=2^2-6\times 2-7\] good so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what do you get when you compute \[g(2)\]?

OpenStudy (anonymous):

2+8=10 and 2^2-12-7=2^2+5 i think?

OpenStudy (anonymous):

oh sorry i confused you. we are ignoring \[f(x)\] for the moment, and just concentrating on \[g(x)\] and we need \[g(2)\]

OpenStudy (anonymous):

what you wrote is correct, \[g(2)=2^2-12-7\] and this number is \[4-12-7=-8-7=-15\] so \[f(2)=-15\]

OpenStudy (anonymous):

now we are not asked for \[f(2)\] we are asked for \[f(g(2))\] and we work from the inside out. we already have computed \[g(2)=-15\] so \[f(g(2))=f(-15)\]

OpenStudy (anonymous):

btw you read "f of g of 2"

OpenStudy (anonymous):

and \[f(x)=x+8\] so \[f(-15)=-15+8=-7\]

OpenStudy (anonymous):

and that is your "final answer" although i sense there might be some confusion as to how we got it

OpenStudy (anonymous):

haha yes just a tad

OpenStudy (anonymous):

ok we can try another one if you like would you like to post one, or would you like me to make one up?

OpenStudy (anonymous):

Let f(x) = 9x – 2 and g(x) = –x + 3. Find f(g(x)).

OpenStudy (amistre64):

another method would be to just insert g(x) into the x-spot of f(x); then make x=2 and solve or in this last case, leave it as is

OpenStudy (anonymous):

ok so once again we work from the inside out. we want "f of g of x"

OpenStudy (amistre64):

f(g(x)) = 9(g(x))-2 replace g(x) with what it equals

OpenStudy (anonymous):

now in this example we do not want \[f(g(2))\] which is a number, but rather we want \[f(g(x))\] which is a function. step one is to replace the general \[g(x)\] which could be any function by the specific one you have, namely \[g(x)=-x+3\] so we write \[f(g(x))=f(-x+3)\]

OpenStudy (anonymous):

so far not much. now comes the only confusing part. and it is confusing because of all the x's you see \[f(x)=9x-2\] but the "x" is not important. it means "multiply by 9, subtract 2"

OpenStudy (amistre64):

it the use of "x" for the independant variable in both cases that makes the notation a bit confusing to begin with

OpenStudy (anonymous):

so we could just as well write \[f(\alpha)=9\alpha -2\] or \[f(\heartsuit)=9\heartsuit -2\]

OpenStudy (anonymous):

and our job now is to replace \[x\] or \[\alpha\]or \[\heartsuit\] by \[-x+3\] (and don't forget parentheses) so you get \[f(-x+3)=9(-x+3)-2\]

OpenStudy (anonymous):

so your answer would be -9x-2??

OpenStudy (anonymous):

then a tiny bit of algebra to clean this up. i will write it in one line, like you would if you were handing it in \[f(g(x))=f(-3x+3)=9(-x+3)-2=-9x+27-2=-9x+25\]

OpenStudy (anonymous):

oooh okayyy

OpenStudy (anonymous):

you have to multiply EVERYTHING in the parentheses by 9

OpenStudy (anonymous):

to recap 1) replace g(x) by the specific on 2) replace the "x" in f(x) by g(x) using parentheses 3) use the distributive property 4) combine like terms the last two steps are algebra steps

OpenStudy (anonymous):

thank you<3